# Compactor: Calibrated Query-Agnostic KV Cache Compression with Approximate Leverage Scores

### v1

@@ 5.5 Ablations
Variant 75% 50% 25% 10% 5%
Full ($\lambda=0.35$) 93.8 87.4 77.5 59.5 43.1
   – Exact $\ell_{i}$ 93.9 88.1 77.5 59.4 43.0
   – $\lambda=0.15$ 93.1 87.8 77.2 60.1 42.1
   – $\lambda=0.2$ 93.2 86.4 77.6 58.6 44.1
   – $\lambda=0.4$ 92.7 88.2 77.8 58.8 42.2
   – No $\vec{a}$ 93.1 81.7 60.2 34.7 24.8
@@ 6 Conclusion
We introduce Compactor , a lightweight, query-agnostic strategy for KV cache pruning that delivers three concrete benefits over prior work: (1) higher accuracy at higher compression, across both synthetic (RULER) and real-world (LongBench) datasets; (2) Compactor scoring incurs minimal overhead compared to the cost of a single Flash Attention [10] pass for large contexts; and (3) Compactor is parameter-free at inference, hardware-agnostic, and orthogonal to other memory-saving techniques. Our context-calibrated compression rule uses only model likelihoods, which makes Compactor an immediately practical drop-in for long-context LLM serving.
@@ Appendix A Proof of Theorem 2
Let data matrix $\mathbf{K}\in\mathbb{R}^{N\times d}$ and target dimension $k=\lceil r\cdot N\rceil$ be given. Let distortion factor $\epsilon\in(0,1)$ and failure probability $\delta\in(0,1)$ be given. Define $\mathbf{\Phi}\in\mathbb{R}^{d\times k}$ be a matrix whose entries are drawn i.i.d from $\mathcal{N}(0,\frac{1}{k})$. Compute the SVD of the sketched matrix $\mathbf{K\Phi}$ and approximate leverage scores ~$\tilde{\ell}_{i}$ as:
$\textsc{svd}\left(\mathbf{K}\mathbf{\Phi}\right)=\mathbf{\tilde{U}}\mathbf{\tilde{\Sigma}}\mathbf{\tilde{V}}^{\top}\qquad\qquad\tilde{\ell}_{i}=||\tilde{U}_{i}||_{2}^{2}$
@@
The proof of the above is straightforward. It relies on the following intermediary result from [14]:
@@
$(1-\epsilon)||U_{i}\mathbf{\Sigma}||^{2}_{2}\leq||\tilde{U}_{i}\tilde{\mathbf{\Sigma}}||_{2}^{2}\leq(1+\epsilon)||U_{i}\mathbf{\Sigma}||^{2}_{2}$
@@ Appendix C Subroutines for Computing Approximate Leverage Scores
Given sketched keys ${\hat{\mathbf{K}}}$, compute the reduced QR factorization ${\hat{\mathbf{K}}}=\mathbf{Q}\mathbf{R}$ where $\mathbf{Q}\in\mathbb{R}^{N\times k}$ has orthonormal columns and $\mathbf{R}\in\mathbb{R}^{k\times k}$ is upper-triangular. The leverage score of the $i$-th token is then
@@
While QR is backward-stable, the orthogonality of $\mathbf{Q}$ degrades when $\kappa({\hat{\mathbf{K}}})$ is large. Re-orthogonalization alleviates this but reduces the potential speed-up.
@@
[…]
@@ Appendix D Subsampled Randomized Hadamard Transform
where $\mathbf{D}$ is a diagonal matrix whose entries are i.i.d. Rademacher (±$\pm 1$) variables, $\mathbf{H}$ is the Walsh–Hadamard matrix (which can be applied with the fast Hadamard transform in time $O(d\log d)$), and $\mathbf{R}$ selects $k$ columns uniformly without replacement. The above construction of $\mathbf{\Phi}$ satisfies the same bounds in Theorem 2 as the Gaussian sketch [2]. The key advantage of utilizing a SRHT is that it can be computed in-place on the KV cache matrices, which can reduce peak memory usage. Empirically we observe that substituting the SRHT for the Gaussian right sketch (all else equal) yields negligible differences in performance on RULER with Llama-3.1-8B. In conjunction with Appendix C, these results further demonstrate that Compactor is relatively robust to implementation details.
@@
Variant 75% 50% 25% 10% 5%
Gaussian 93.8 87.4 77.5 59.5 43.1
SRHT 93.7 87.8 77.4 59.9 42.4
Appendix E Longbench Code Completion
As mentioned in the main body of the paper, we exclude Longbench code-completion tasks from evaluation with other query-agnostic tasks because of the lack of a defined query. For completeness, we include results on the code-portion of Longbench in Table 5. We observe that Compactor achieves close to baseline performance on Llama 3.1, despite being evaluated in a query-aware setting. SnapKV demonstrates the best performance by a small margin, showing the utility of the recent token window for code completion tasks. Such an observation is not surprising: the code to be completed disproportionately relies on the last window of text to determine what exactly will be completed. The increase in performance of all methods under compression for Qwen 2.5 suggests that the entire context is not needed for the task at hand.
Table 5: Longbench scores for code completion. Cells in blue denote our contribution.
Task Method Llama 3.1 Qwen 2.5
100% 50% 25% 10% 5% 100% 50% 25% 10% 5%
Code Completion Compactor 0.253 0.250 0.248 0.231 0.229 0.255 0.263 0.281 0.292 0.244
SnapKV 0.255 0.249 0.256 0.223 0.275 0.290 0.288 0.259
PyramidKV 0.247 0.254 0.224 0.224 0.270 0.289 0.272 0.233

### v2

@@ 5.5 Ablations
Variant 75% 50% 25% 10%
Full 95.6 94.7 83.1 64.8
   – Exact $\ell_{i}$ 95.5 94.9 82.8 64.7
   – $k=64$ 95.3 94.2 83.2 64.8
   – $\lambda=0.4$ 95.7 94.8 82.3 63.8
   – $\lambda=0.2$ 94.1 94.3 83.1 63.0
   – $\lambda=0.15$ 94.3 94.5 80.1 62.3
   – Only $\vec{a}$ 93.5 79.0 61.3 35.7
   – Only $\vec{o}$ 96.1 94.1 73.5 44.7
@@ 6 Conclusion
We introduce Compactor , a lightweight, query-agnostic strategy for KV cache pruning that delivers three concrete benefits over prior work: (1) higher accuracy at higher compression, across both synthetic (RULER) and real-world (LongBench) datasets; (2) Compactor scoring incurs minimal overhead compared to the cost of a single Flash Attention (dao_flashattention_2022) pass for large contexts; and (3) Compactor is training-free at inference and orthogonal to other memory-saving techniques. Finally, we are the first to introduce context-calibrated compression for automatic compression ratio selection, which is applicable to any compression method. We believe Compactor is an immediately practical drop-in for long-context LLM serving.
@@ Appendix A Proof of Theorem 2
Let data matrix $\mathbf{K}\in\mathbb{R}^{N\times d}$ and target dimension $k=\lceil r\cdot N\rceil$ be given. Let distortion factor $\epsilon\in(0,1)$ and failure probability $\delta\in(0,1)$ be given. Define $\mathbf{\Phi}\in\mathbb{R}^{d\times k}$ be a matrix whose entries are drawn i.i.d from $\mathcal{N}(0,\frac{1}{k})$. Compute the SVD of the sketched matrix $\mathbf{K\Phi}$ and approximate leverage scores $\tilde{\ell}_{i}$ as:
[…]
@@
The proof of the above is straightforward. It relies on the following intermediary result from (gilbert_sketched_2012):
@@
[…]
@@ Appendix C Subroutines for Computing Approximate Leverage Scores
Given sketched keys $^{\hat{\mathbf{K}}}$, compute the reduced QR factorization ${\hat{\mathbf{K}}}=\mathbf{Q}\mathbf{R}$ where $\mathbf{Q}\in\mathbb{R}^{N\times k}$ has orthonormal columns and $\mathbf{R}\in\mathbb{R}^{k\times k}$ is upper-triangular. The leverage score of the $i$-th token is then
$\tilde{\ell}_{i}\;=\;\bigl\lVert\mathbf{Q}_{i,:}\bigr\rVert_{2}^{\,2},\quad i=1,\dots,N.$
@@
and set $\tilde{\ell}_{i}=\lVert\widetilde{\mathbf{U}}_{i,:}\rVert_{2}^{\,2}$. The eigendecomposition is the most sensitive of the three routes. Small singular values of $^{\hat{\mathbf{K}}}$ are squared and may fall below machine precision, inflating the corresponding $\mathbf{\Lambda}^{-1/2}$ entries and corrupting $\tilde{\mathbf{U}}$. Clamping $\lambda_{j}\leftarrow\max(\lambda_{j},\varepsilon)$ with $\varepsilon\!\approx\!10^{-6}$ largely mitigates the issue.
@@ Appendix D Subsampled Randomized Hadamard Transform
where $\mathbf{D}$ is a diagonal matrix whose entries are i.i.d. Rademacher ($\pm 1$) variables, $\mathbf{H}$ is the Walsh–Hadamard matrix (which can be applied with the fast Hadamard transform in time $O(d\log d)$), and $\mathbf{R}$ selects $k$ columns uniformly without replacement. The above construction of $\mathbf{\Phi}$ satisfies the same bounds in Theorem 2 as the Gaussian sketch (ailon_fast_2009). The key advantage of utilizing a SRHT is that it can be computed in-place on the KV cache matrices, which can reduce peak memory usage. Empirically we observe that substituting the SRHT for the Gaussian right sketch (all else equal) yields negligible differences in performance on RULER with Llama-3.1-8B. In conjunction with Appendix C, these results further demonstrate that Compactor is relatively robust to implementation details.
@@
Variant 75% 50% 25% 10%
Gaussian 95.6 94.7 83.1 64.8
SRHT 95.4 94.3 84.1 64.0